# Concept-skill Transferability-based Data Selection for Large Vision-Language Models

### v1

@@ Abstract
Instruction tuning, or supervised finetuning on extensive task-specific data, is necessary for Large Vision-Language Models (LVLMs) to generalize well across a broad range of vision-language (VL) tasks. However, training on large VL datasets can become prohibitively expensive. In this work, we introduce COINCIDE, an effective and scalable data selection technique that uses a small model as a reference model to select visual instruction tuning data for efficient finetuning of a target LVLM, focusing on diversity and transferability. Specifically, we cluster the training data using internal activations from a small model, which identifies VL concept-skill compositions needed by a target LVLM. We then sample data from these diverse clusters by considering their density and transferability, or the ability to transfer well to other concept-skill compositions. This approach ensures the diversity of these compositions, which is vital for LVLM generalization. Extensive experiments demonstrate that COINCIDE achieves superior performance and data selection efficiency against 8 strong baselines on two distinct datasets: LLaVA-1.5 and Vision-Flan. Using only 20% of the LLaVA-1.5 dataset, COINCIDE achieves performance comparable to the LVLM finetuned on the whole dataset, with 70% reduction of the wall-clock running time. On the Vision-Flan dataset, our method achieves superior results with only 16.7% of the training data. $\dagger$ Equal advising
@@
Jaewoo Lee1   Boyang Li†,2   Sung Ju Hwang†,1,3 KAIST1 Nanyang Technological University, Singapore2, DeepAuto3 jwlee8877@kaist.ac.kr boyang.li@ntu.edu.sg sjhwang82@kaist.ac.kr
@@ 1 Introduction
In this paper, we investigate the selection of a coreset, a subset that approximates the performance of the full dataset, from large VIT datasets. Conventional coreset selection approaches (Marion et al., 2023; Zhou et al., 2023; Chen et al., 2023a) usually measure training data quality with a score metric to select valuable training data. However, we discover a mismatch between these score metrics and the highly diverse nature of VIT datasets. Due to the divergence within VIT datasets, selecting any subset based on a single metric leads to a coreset dominated by a few tasks. As shown in Figure 1, the coreset from the middle of EL2N (Paul et al., 2021) score distribution consists of samples mostly from only 3-4 tasks. This bias severely reduces the diversity of the selected coreset and, in our experiments, weakens LVLM generalization (Table 1).
Figure 2: VL tasks (e.g., VQAv2 and GQA, LLaVA-Conv and LLaVA-Reason) share VL concept-skill compositions.
Instead of coreset selection at the dataset level, which involves applying a single score metric across the entire dataset, we propose to select data at the level of data clusters, which roughly corresponds to compositions of VL concepts and skills. For example, a concept could be street signs or trains on a railroad, while a skill could be OCR, recognizing color, or reasoning. Upon close inspection, we find that different VL tasks contain overlap over these concept-skill compositions. As exemplified in Figure 2, LLaVA-Conv and LLaVA-Reason contain questions about the risks of snowboard jumps, despite their separate focuses on multi-turn conversations and reasoning. This suggests sampling over the clusters would be more effective in enhancing the diversity of VL concept-skill compositions than sampling over datasets or tasks.
To this end, we introduce COre INstruction Concept-skIll Data Election (COINCIDE), which identifies VL concept-skill compositions through data clustering using activations from an off-the-shelf, small VLM (Figure 3 Left). From each cluster, COINCIDE selects training data for a target LVLM by considering transferability (i.e., how well knowledge from each cluster can facilitate LVLM’s learning in other clusters) and internal density of clusters (Figure 3 Right). Empirically, we find that transferability correlates well with cosine similarity among clusters. Based on the findings, we select more data points from more transferable clusters. Further, we sample fewer data points from denser clusters, as data points in dense clusters are likely redundant. By selecting data from diverse clusters, COINCIDE enhances the diversity of VL concept-skill compositions in the selected data, leading to better LVLM generalization.
@@ 2 Related Work
Figure 3: Illustration of COINCIDE. Our method utilizes a small VLM to cluster visual instruction tuning data based on concept-skill compositions. We then assess the cluster transferability as the mean cosine similarity to other cluster centroids. We further compute the cluster density as the mean Gaussian kernel distance among all data pairs within the cluster. Using cluster transferability and density, COINCIDE determines the number of data to sample from each cluster and performs intra-cluster sampling. Finally, it combines all the selected samples from all the clusters to compose the final coreset.
@@
Discovering concepts learned by neural networks is a popular topic in interpretability research (Kim et al. 2018; FEL et al. 2023; Manning et al. 2020). Notably, Kowal et al. (2024) performs hierarchical clustering in layer-wise activation space. Tiong et al. (2024) attempt to identify latent skills underlying VL datasets. Michaud et al. (2023) performs spectral clustering to discover the skills of LLMs. Though these works provide inspiration, they are orthogonal to our work, whose main objective is to sample from data clusters rather than understanding existing neural networks. The only application of concept discovery we are aware of is by Gupta et al. (2017), who shows enforcing consistent VL concepts improves transfer learning.
@@ 3 Method
We start by introducing the framework that utilizes neuron activations from a small VLM to group VIT data into clusters, where each cluster comprises samples exhibiting a similar concept-skill composition (Section 3.2). Next, we conduct experiments to examine the correlation between the similarity of a cluster centroid to other centroids and the transferability of that cluster to others (Section 3.3). Based on our findings, we describe our data selection strategy, which performs cluster-wise sample selection by selecting different numbers of samples from clusters depending on their transferability and diversity (Section 3.4). The overall framework of our approach is illustrated in Figure 3.
@@ 3.2 Discovering Concept-Skill Compositions
We cluster VIT training data points using their feature vector from multiple layers of a small VLM, called a reference model. We extract the features right after the MSA of the $l$-th layer (Eq. 1) and process them into unit-length vectors:
@@
where the mean-pooling is performed across the number of visual and text tokens, respectively. The hyperbolic tangent function, tanh, is necessary to reduce the impact of a few extreme activations, which are described by Sun et al. (2024). Without this step, these large values would dominate the feature vector and skew the clustering. After that, we concatenate features from the small VLM’s layers:
@@
Then, we perform spherical k-means clustering on $\bm{u}^{m}$, yielding $K$ clusters. To ensure the purity of clusters, we set $K$ to a large number, such as $10,000$. Despite its simplicity, the k-means procedure runs in $O(NK)$ time for $N$ data points, which is advantageous when both $N$ and $K$ are large. Other clustering techniques such as spectral clustering or affinity propagation are much more expensive. Qualitative analysis indicates the clusters effectively capture concept-skill compositions. We provide visualization of the clusters in Appendix C.
@@ 4.1 Setup
For the target LVLMs, we use the pre-trained LLaVA-1.5 model (Liu et al., 2023a) with a default size of 7B parameters unless otherwise specified. In all experiments, we train the models using LoRA (Hu et al., 2022) for one epoch, following the official finetuning hyperparameters specified in LLaVA-1.5. As a reference model, we use the TinyLLaVA-2B (Zhou et al., 2024), a small VLM finetuned on the target VIT dataset, for efficient coreset selection for all methods unless otherwise specified. All experiments are conducted using 4 V100 GPUs.
@@ 4.2 Results and Discussion
Figure 5: Average relative performances of all techniques at different coreset sizes for the LLaVA-1.5 dataset.
@@
Another curious phenomenon is that several baselines, including CLIP-Score, Perplexity, and Self-Filter, experience performance declines as the sampling ratio increases in Figure 6. A similar trend is observed in the random baseline in Figure 5. This underscores the importance of coreset selection, as merely increasing the dataset size does not guarantee improved LVLM capabilities.
Figure 6: Average relative performances of all techniques at different coreset sizes for the Vision-Flan dataset.
@@
In Figure 7, we plot the wall-clock time cost of the entire pipeline of data selection and model finetuning versus the average relative performance (Rel.) on the LLaVA-1.5 dataset. COINCIDE achieves 97.4%, 98.4%, and 99.4% Rel. with the wall-clock times of 15.1, 25.1, and 35.1 hours, respectively. In contrast, finetuning on all data takes 50 hours.
@@ 5 Conclusion
In this paper, we introduce COINCIDE, a cluster-level visual instruction tuning data selection for efficient finetuning of Large Vision-Language Models. We demonstrate that clustering based on inner activations from a small model can represent visual-linguistic concept-skill compositions shared among diverse tasks in visual instruction tuning datasets. Additionally, our empirical investigation validates a strong positive correlation between cosine similarity and transferability among clusters. Based on the transferability and density of clusters, COINCIDE selects more samples from more transferable and less dense clusters to enhance training efficacy, while preserving the diversity of concept-skill compositions within the coreset to ensure better model generalization ability. Comprehensive experiments on the LLaVA-1.5 and Vision-Flan datasets demonstrate that our method outperforms baselines across several benchmarks with the lowest data selection cost, showcasing its effectiveness and efficiency.
@@
Additionally, we assess whether the data with similar concept-skill compositions are concentrated well on the clusters through human inspection. Therefore, further investigation should be conducted to quantitatively evaluate the clustering of data with similar concept-skill compositions. Quantitative measures are expected to enable more accurate identification of VL concept-skill compositions and their transferability.
@@
In this work, we use publicly available visual instruction tuning datasets for coreset selection to enable easy replication. However, some data in the datasets contain erroneous answers about the visual content or images that do not clearly connect with the provided answers. Finetuning Large Vision-Language Models (LVLMs) with such data conveys wrong interpretations, inducing hallucinations in the LVLMs. Hallucination in LVLMs refers to a phenomenon where the LVLMs generate descriptions that are inconsistent with the target image. This poses a significant ethical issue for deploying LVLM in real-world applications. However, current coreset selection techniques, including ours, do not address hallucination in their selection processes. This motivates further research in coreset selection to identify visual instruction tuning data that minimizes hallucinations, aiming to build more reliable and trustworthy LVLMs.
@@ Appendix A Details of Experimental Setups
    Figure 8: Task-wise transferability. We group the VIT data by task names and average the cluster transferability of each data.
@@ Appendix B Visualizing LVLM Skills with Relevancy Maps
For the investigation, we inspect the 4th, 8th, 12th, 16th, and 20th layers of the TinyLLaVA-2B (Zhou et al., 2024) model and identify the layer that activates the most relevant visual regions. The results, shown in Figure 10, reveal that (1) the most relevant layer varies according to the concept-skill composition and (2) the most relevant layer is the same across diverse VIT data when the data shares a similar concept-skill composition. These findings support our initial assumption that different layers contribute to distinct concepts and skills. Therefore, using neuron activations from diverse layers can effectively group VIT data according to their concept-skill composition.
@@ Appendix D In-Depth Analysis on Concept-Skill Composition Transferability
Table 5: Transfering to the larger target model. We validate if the coresets selected from TinyLLaVA-2B are transferable to LLaVA-1.5-13B finetining. We train the LLaVA-1.5-13B using coresets with 20% sampling ratio and estimate performance on various multimodal benchmarks. The best and the second best results are highlighted in bold and underline, respectively.
@@ F.1 Transfering to Larger Target Model
We evaluate the performance of the larger target model (LLaVA-1.5-13B) finetuned on coresets gathered by the small VLM (TinyLLaVA-2B). Table 5 summarizes the performances across various benchmarks. The results demonstrate the effectiveness of our method in selecting a coreset that can be successfully transferred to the larger target model.
Table 7: We investigate the impact of multimodal neuron activations.
Neuron Activation Rel. (%)
Boolean 95.7
Last layer 96.5
MSA layers 97.4
FFN layers 96.0
@@ F.3 Hyperparameters
We conduct ablation studies on hyperparameters of our method, which include the number of clusters ($\bm{K}$) and the temperature ($\bm{\tau}$). The results, summarized in Section F.1, reveal that a sufficiently large number of clusters is essential to ensure cluster purity and diversity of VL concept-skill compositions, ensuring effective representation of the compositions and enhancing the generalization ability of LVLM. Furthermore, we find that setting the temperature too low leads to a biased coreset selection, as most samples are then selected from a few clusters. This undermines the diversity within the coreset, leading to a decline in overall performance.
@@ F.4 Multimodal Neuron Activation
We further analyze the impact of different multimodal neuron activations on the performance of our method. COINCIDE selects neuron activations from the MSA blocks across the 4th, 8th, 12th, 16th, and 20th layers of the reference model. We experiment with different neuron activations and present the results in Section F.1. Transforming the neuron activations from the MSA blocks into boolean vectors by mapping negative values to -1 and positive values to 1 causes a significant performance drop, likely due to substantial information loss, yielding inaccurate clustering and transferability calculation. Extracting neuron activations only from the last layer of the reference model causes a slight performance decrease. As discussed in Section 3.2, LVLM abilities stem from various layers. Hence, relying on the last layer captures only a small portion of these capabilities, leading to the performance decline. Finally, utilizing the neuron activations from the MSA blocks gives superior performance compared to using activations from the FFN blocks. We believe this is because MSA layers use self-attention to share multimodal information, providing richer multimodal understanding.

### v2

@@ Abstract
Instruction tuning, or supervised finetuning on extensive task-specific data, is necessary for Large Vision-Language Models (LVLMs) to generalize well across a broad range of vision-language (VL) tasks. However, training on large VL datasets can become prohibitively expensive. In this work, we introduce COINCIDE, an effective and scalable data selection technique that uses a small model as a reference model to select visual instruction tuning data for efficient finetuning of a target LVLM, focusing on diversity and transferability. Specifically, we cluster the training data using internal activations from a small model, which identifies VL concept-skill compositions needed by a target LVLM. We then sample data from these diverse clusters by considering their density and transferability, or the ability to transfer well to other concept-skill compositions. This approach ensures the diversity of these compositions, which is vital for LVLM generalization. Extensive experiments demonstrate that COINCIDE achieves superior performance and data selection efficiency against 8 strong baselines on two distinct datasets: LLaVA-1.5 and Vision-Flan. Using only 20% of the LLaVA-1.5 dataset, COINCIDE achieves performance comparable to the LVLM finetuned on the whole dataset, with 70% reduction of the wall-clock running time. On the Vision-Flan dataset, our method achieves superior results with only 16.7% of the training data. Our code is available at https://github.com/G-JWLee/COINCIDE_code. $\dagger$ Equal advising
@@
Jaewoo Lee1   Boyang Li†,2   Sung Ju Hwang†,1,3 KAIST1 Nanyang Technological University, Singapore2 DeepAuto3 jwlee8877@gmail.com boyang.li@ntu.edu.sg sjhwang82@kaist.ac.kr
@@ 1 Introduction
In this paper, we investigate the selection of a coreset, a subset that approximates the performance of the full dataset, from large VIT datasets. Conventional coreset selection approaches (Marion et al., 2023; Zhou et al., 2023; Chen et al., 2023a) usually utilize a score metric to select training data. As VIT datasets are highly diverse and feature multiple data modes (Figure 1), data selection using any single metric would produce a coreset dominated by a few tasks. Figure 1 indicates that, selecting 20% of data from any part of the metric distribution of EL2N (Paul et al., 2021) or Self-Filter (Chen et al., 2024a) would exclude many data modes, which severely reduces the diversity of the selected coreset and harms generalization. As our experiments show (Table 1), this type of coreset selection degrades LVLM performance.
Figure 2: Different VL tasks (e.g., VQAv2 and GQA, LLaVA-Conv and LLaVA-Reason) share VL concept-skill compositions.
Our solution to the multitude of data modes is straightforward: we explicitly identify the modes by clustering the VIT data points using features from multiple layers in a small LVLM. Interestingly, we observe that the clusters thus identified roughly coincide with compositions of VL concepts and skills. For example, a concept could be street signs or trains on a railroad, while a skill could be OCR, recognizing color, or reasoning. Upon close inspection, we find that different VL tasks contain overlap over these concept-skill compositions. As exemplified in Figure 2, LLaVA-Conv and LLaVA-Reason contain questions about the risks of snowboard jumps, despite their separate focuses on multi-turn conversations and reasoning. This suggests sampling over the clusters would be more effective in enhancing the diversity of VL concept-skill compositions than sampling over datasets or tasks.
To this end, we introduce COre INstruction Concept-skIll Data Election (COINCIDE), which identifies VL concept-skill compositions through data clustering using activations from an off-the-shelf, small LVLM (Figure 3 Left). From each cluster, COINCIDE selects training data for a target LVLM by considering transferability (i.e., how well knowledge from each cluster can facilitate LVLM’s learning in other clusters) and internal density of clusters (Figure 3 Right). Empirically, we find that transferability correlates well with cosine similarity among clusters. Based on the findings, we select more data points from more transferable clusters. Further, we sample fewer data points from denser clusters, as data points in dense clusters are likely redundant.
@@ 2 Related Work
Figure 3: Illustration of COINCIDE. Our method utilizes a small LVLM to cluster visual instruction tuning data based on concept-skill compositions. We then assess the cluster transferability as the mean cosine similarity to other cluster centroids. We further compute the cluster density as the mean Gaussian kernel distance among all data pairs within the cluster. Using cluster transferability and density, COINCIDE determines the number of data to sample from each cluster and performs intra-cluster sampling. Finally, it combines all the selected samples from all the clusters to compose the final coreset.
@@
Concept discovery in neural networks is a key topic in interpretability research (Kim et al. 2018; FEL et al. 2023; Manning et al. 2020). Notably, Kowal et al. (2024) performs hierarchical clustering in layer-wise activation space. Tiong et al. (2024) attempts to identify latent skills underlying VL datasets. Michaud et al. (2023) performs spectral clustering to discover LLMs skills. Though these works provide inspiration, they are orthogonal to our work, whose main objective is to sample from data clusters rather than understanding existing neural networks. The only application of concept discovery we are aware of is by Gupta et al. (2017), showing consistent VL concepts improve transfer learning.
@@ 3 Method
We start by introducing the framework that utilizes neuron activations from a small LVLM to group VIT data into clusters, where each cluster comprises samples exhibiting a similar concept-skill composition (Section 3.2). Next, we conduct experiments to examine the correlation between the similarity of a cluster centroid to other centroids and the transferability of that cluster to others (Section 3.3). Based on our findings, we describe our data selection strategy, which performs cluster-wise sample selection by selecting different numbers of samples from clusters depending on their transferability and diversity (Section 3.4). The overall framework of our approach is illustrated in Figure 3.
@@ 3.2 Discovering Concept-Skill Compositions
We cluster VIT training data points using their feature vector from multiple layers of a small LVLM, called a reference model. We extract the features right after the MSA of the $l$-th layer (Eq. 1) and process them into unit-length vectors:
@@
where the mean-pooling is performed across the number of visual and text tokens, respectively. The hyperbolic tangent function, tanh, is necessary to reduce the impact of a few extreme activations, which are described by Sun et al. (2024). Without this step, these large values would dominate the feature vector and skew the clustering. After that, we concatenate features from the small LVLM’s layers:
@@
Then, we perform spherical k-means clustering on $\bm{u}^{m}$, yielding $K$ clusters. To ensure the purity of clusters, we set $K$ to a large number, such as $10,000$. Despite its simplicity, the k-means procedure runs in $O(NK)$ time for $N$ data points, which is advantageous when both $N$ and $K$ are large. Other clustering techniques such as spectral clustering or affinity propagation are much more expensive. Qualitative analysis indicates the clusters coincide with concept-skill compositions. We provide visualization of the clusters in Appendix C.
@@ 4.1 Setup
For the target LVLMs, we use the pre-trained LLaVA-1.5 model (Liu et al., 2023a) with a default size of 7B parameters unless otherwise specified. In all experiments, we train the models using LoRA (Hu et al., 2022) for one epoch, following the official finetuning hyperparameters specified in LLaVA-1.5. As a reference model, we use the TinyLLaVA-2B (Zhou et al., 2024), a small LVLM finetuned on the target VIT dataset, for efficient coreset selection for all methods unless otherwise specified. All experiments are conducted using 4 V100 GPUs.
@@ 4.2 Results and Discussion
Figure 5: Average relative performances of all coreset selection techniques at different sampling ratios for the LLaVA-1.5 dataset.
@@
Another curious phenomenon is that several baselines, including CLIP-Score, Perplexity, and Self-Filter, experience performance declines as the sampling ratio increases in Figure 6. A similar trend is observed in the random baseline in Figure 5. This underscores the importance of deliberate coreset selection, as merely increasing the dataset size does not guarantee improved LVLM capabilities.
Figure 6: Average relative performances of all coreset selection techniques at different sampling ratios for the Vision-Flan dataset.
@@
In Figure 7, we plot the wall-clock time cost of the entire pipeline of data selection and model finetuning versus the average relative performance (Rel.) on the LLaVA-1.5 dataset. COINCIDE achieves 97.4%, 98.4%, and 99.4% relative performance with the wall-clock times of 15.1, 25.1, and 35.1 hours, respectively. In contrast, finetuning on all data takes 50 hours.
@@ 5 Conclusion
In this paper, we introduce COINCIDE, a cluster-level data selection technique for efficient visual instruction tuning of Large Vision-Language Models. We demonstrate that clustering based on internal activations from a small model can represent visual-linguistic concept-skill compositions shared among diverse tasks in visual instruction tuning datasets. Additionally, our empirical investigation validates a strong positive correlation between cosine similarity and transferability among clusters. Based on the transferability and density of clusters, COINCIDE selects more samples from more transferable and less dense clusters to enhance training efficacy, while preserving the diversity of concept-skill compositions within the coreset to ensure better model generalization ability. Comprehensive experiments on the LLaVA-1.5 and Vision-Flan datasets demonstrate that our method outperforms baselines across several benchmarks with the lowest data selection cost, showcasing its effectiveness and efficiency. The success of COINCIDE suggests redundancy in popular VIT datasets and underscores the importance of a thorough understanding of data in training LVLMs.
@@
Additionally, we assess whether the data with similar concept-skill compositions are concentrated well on the clusters through human inspection. Therefore, further investigation should be conducted to quantitatively evaluate the clustering of data with similar concept-skill compositions, which may enable accurate identification of VL concept-skill compositions and accurate quantification of their transferability.
@@
In this work, we use publicly available visual instruction tuning datasets for coreset selection to enable easy replication. However, some data in the datasets contain erroneous answers about the visual content or images that do not clearly connect with the provided answers. Finetuning Large Vision-Language Models (LVLMs) with such data may lead to the generation of erroneous interpretations of images or hallucinations. This may pose an ethical issue for LVLM deployment in the real world. However, current coreset selection techniques, including ours, do not address hallucination in their selection processes. This motivates further research in coreset selection to identify visual instruction tuning data that minimizes hallucinations, aiming to build more reliable and trustworthy LVLMs.
Acknowledgements
Jaewoo Lee and Sung Ju Hwang are supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00256259) and a grant of the Korea Machine Learning Ledger Orchestration for Drug Discovery Project (K-MELLODDY), funded by the Ministry of Health & Welfare and Ministry of Science and ICT, Republic of Korea (No. RS-2024-12345678). Boyang Li is supported by the Nanyang Associate Professorship and Fellowship (NRF-NRFF13-2021-0006) of the National Research Foundation, Singapore. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the funding agencies.
@@ Appendix A Details of Experimental Setups
    Figure 8: Task-wise transferability. We group the VIT data based on task names and then report the average cluster transferability of each group.
@@ Appendix B Visualizing LVLM Skills with Relevancy Maps
For the investigation, we inspect the 4th, 8th, 12th, 16th, and 20th layers of the TinyLLaVA-2B (Zhou et al., 2024) model and identify the layer that activates the most relevant visual regions. The findings in Figure 10 reveal that (1) the most relevant layer varies according to the concept-skill composition and (2) the most relevant layer is the same across diverse VIT data when the data shares a similar concept-skill composition. This supports our assumption that different layers contribute to distinct concepts and skills, allowing neuron activations from various layers to effectively group VIT data by their concept-skill composition.
@@ Appendix D In-Depth Analysis on Concept-Skill Composition Transferability
Table 5: Transferring to the larger target model. We validate if the coresets selected from TinyLLaVA-2B are transferable to LLaVA-1.5-13B finetining. We train the LLaVA-1.5-13B using coresets with 20% sampling ratio and estimate performance on various multimodal benchmarks. The best and the second best results are highlighted in bold and underline, respectively.
@@ Appendix E Concept-Skill Diversity within Coresets
Table 6: We investigate the impact of various representations of multimodal neuron activation.
Neuron Activation Rel. (%)
Boolean 95.7
Last layer 96.5
MSA layers 97.4
FFN layers 96.0
@@ F.1 Transfering to Larger Target Model
We evaluate the performance of the larger target model (LLaVA-1.5-13B) finetuned on coresets gathered by the small LVLM (TinyLLaVA-2B). Table 5 summarizes the performances across various benchmarks. The results demonstrate the effectiveness of our method in selecting a coreset that can be successfully transferred to the larger target model.
@@ F.3 Hyperparameters
We conduct ablation studies on hyperparameters of our method, which include the number of clusters ($\bm{K}$) and the temperature ($\bm{\tau}$). The results, summarized in Appendix E, reveal that a sufficiently large number of clusters is essential to ensure cluster purity and diversity of VL concept-skill compositions, ensuring effective representation of the compositions and enhancing the generalization ability of LVLM. Furthermore, we find that setting the temperature too low leads to a biased coreset selection, as most samples are then selected from a few clusters. This undermines the diversity within the coreset, leading to a decline in overall performance.
@@ F.4 Multimodal Neuron Activation
We further analyze the impact of different multimodal neuron activations on the performance of our method. COINCIDE selects neuron activations from the MSA blocks across the 4th, 8th, 12th, 16th, and 20th layers of the reference model. We experiment with different neuron activations and present the results in Appendix E. Transforming the neuron activations from the MSA blocks into boolean vectors by mapping negative values to -1 and positive values to 1 causes a significant performance drop, likely due to substantial information loss, yielding inaccurate clustering and transferability calculation. Extracting neuron activations only from the last layer of the reference model causes a slight performance decrease. As discussed in Section 3.2, LVLM abilities stem from various layers. Hence, relying on the last layer captures only a small portion of these capabilities, leading to the performance decline. Finally, utilizing the neuron activations from the MSA blocks gives superior performance compared to using activations from the FFN blocks. We believe this is because MSA layers use self-attention to share multimodal information, providing richer multimodal understanding.